Calculate the mass of water vapor produced when 6.00 L of propane gas undergo complete combustion at a temperature of 250°C and a pressure of 1.00 atm.
1
Expert's answer
2017-04-17T09:28:05-0400
C_3 H_8+〖5O〗_2→〖3CO〗_2+〖4H〗_2 O
PV=nRT ⇒n= PV/RT
n(C_3 H_8)= (1*101,3*6.00)/(8.31*(250+273))=0,14 (mol)
n(H_2 O)=4*n(C_3 H_8 )= 4*0.14= 0.56 (mol)
m(H_2 O)=n(H_2 O)*M(H_2 O)=0.56*18=10.07 (g)

Solution: 10.07 g.